### v1

# Kinematics of Circumgalactic OĀ VIĀ Gas and Disk Rotation of Star-forming Galaxies

Abstract
Quasar sightline observations reveal that low-ionization-state gas corotates with the galaxy disk and often at sub-centrifugal velocities, suggesting that the gas is spiraling towards the galaxy disk. However, while observations ubiquitously detect OĀ VIĀ absorption around low-redshift, Ā star-forming galaxies, the relationship between OĀ VIĀ and the galaxy disk, especially the kinematics, is not well-established. This work focuses on the OĀ VIĀ kinematics and its comparison with that of the low ions and galactic disk rotation. We present observations of 18 pairs of quasars and star-forming galaxies. All quasar sightlines intersect the circumgalactic medium (CGM) within 45āĀ from the galaxy major axes. We show that while individual OĀ VIĀ velocity components do not correlate with disk rotation, the bulk of OĀ VIĀ gas in individual sightlines rarely counter-rotates. We then match OĀ VIĀ velocity components with those of low ions by minimizing the difference of their velocity centroids. The OĀ VIĀ velocity components with successful low-ion matches are typically found at small sightline impact parameters and are more likely to corotate with the disk. We suggest that the low-ion-matched OĀ VIĀ velocity components trace the gas co-spatial with the low ions near the extended disk plane in the inner CGM, whereas those without low-ion matches represent the gas at large 3D radii. While the gas at large radii is theoretically expected to kinematically correlate with the disk angular momentum, this correlation is expected to be weaker due to the higher turbulent to mean rotation velocity ration at large radii, consistent with our results.
1 Introduction
The interplay between galaxies and their surrounding gas shapes the evolution of galaxies. Galaxies gather gas from their surroundings to form stars, whereas winds driven by supernovae and active galactic nuclei (AGN) remove gas from the galaxies and slow star formation. The surrounding gas, known as the circumgalactic medium (CGM), is a substantial reservoir of baryons and metals and extends at least to the virial radius () of galaxies (see Tumlinson etĀ al. 2017 and Faucher-GiguĆØre & Oh 2023 for reviews). Because the low gas density makes direct imaging of the CGM challenging, quasar spectroscopy serves as a powerful tool for detecting the CGM in absorption and characterizing CGM properties.
The ubiquitous detection of OĀ VI 1031, 1037Ā absorption around low-redshift, Ā star-forming galaxies in contrast to the low detection rate around quiescent galaxies have received significant attention (Tumlinson etĀ al., 2011; Tchernyshyov etĀ al., 2023). Dwarf galaxies also show the same contrast (Dutta etĀ al., 2024), even though the OĀ VIĀ detection rate and column density of star-forming dwarf galaxies are lower than that of Ā galaxies (Johnson etĀ al., 2015, 2017; Tchernyshyov etĀ al., 2022; Mishra etĀ al., 2024). The origin of this āOĀ VIĀ dichotomyā remains controversial. AGN feedback possibly leads to a low CGM mass fraction, resulting in the low OĀ VIĀ detection rate of quiescent galaxies (Suresh etĀ al., 2017; Nelson etĀ al., 2018; Davies etĀ al., 2020). Some work argued that the quiescent galaxies have higher average halo masses than the star-forming galaxies in the observed galaxy samples, and hence, the higher virial temperatures promote oxygen to higher ionization states. In particular, the predicted CGM properties change significantly for galaxies with halo masses around Ā due to virialization (Oppenheimer etĀ al., 2016; Mathews & Prochaska, 2017; Stern etĀ al., 2018; Sultan etĀ al., 2025). However, recent work with a mass-controlled galaxy sample still found the āOĀ VIĀ dichotomyā. This implies a potential connection between OĀ VIĀ and ongoing star formation, for which the transformation of CGM happens concurrently with the quenching of galaxies (Tchernyshyov etĀ al., 2023).
Dedicated efforts with quasar absorption-line studies have analyzed the circumgalactic O VI properties of star-forming galaxies and compared that with the cool gas ( K) traced by low-ionization-state ions (e.g., Mg II). For  star-forming galaxies, the cool gas is typically detected only in the inner CGM of kpc, and the strength of the low-ionization-state absorption decreases sharply with increasing projected separation from the galaxies (Chen et al., 2010; Nielsen et al., 2013; Huang et al., 2021). In contrast, O VI is detected in sightlines out to at least two to three times the galaxy virial radius, and the O VI column density only declines slowly with the sightline impact parameters (Tumlinson et al., 2011; Tchernyshyov et al., 2022; Qu et al., 2024). Similarly for star-forming dwarf galaxies, O VI has a higher detection rate and a more extended distribution compared to that of the low ions (Mishra et al., 2024; Dutta et al., 2024). But surprisingly, both low ions and O VI exhibit a bimodal distribution with azimuthal angle111 The azimuthal angle is defined as the angle between the galaxy major axis and the line drawn between the center of the galaxy and the quasar. (Bouché et al., 2012; Kacprzak et al., 2012, 2015; Schroetter et al., 2019). The bimodality in gas distribution is also supported by recent cloud-by-cloud multiphase, ionization modelings that indicate a larger number of clouds in sightlines near the galaxy major and minor axes (Sameer et al., 2024).
In addition to the spatial distribution, kinematics provides crucial information for understanding the origin and properties of the CGM traced by OĀ VIĀ and low-ionization-state ions. While OĀ VIĀ absorption has a larger velocity spread compared to that of the low ions, they typically span similar velocity ranges (Werk etĀ al., 2016; Nielsen etĀ al., 2017). However, while the absorption of low ions becomes broader and stronger towards the galaxy minor axes, a result attributed to gas kinematically disturbed by outflows (Bordoloi etĀ al., 2011; Martin etĀ al., 2019; Schroetter etĀ al., 2019), the OĀ VIĀ velocity spread does not vary with the sightline azimuthal angle and galaxy disk inclination (Nielsen etĀ al., 2017). Furthermore, for sightlines near the galaxy major axes, the cool gas (traced by MgĀ II) corotates with the galaxy disk and often at sub-centrifugal velocities (Steidel etĀ al., 2002; Kacprzak etĀ al., 2010, 2011; Ho etĀ al., 2017; Zabl etĀ al., 2019), suggesting that the gas is infalling (Ho & Martin, 2020). On the other hand, the lack of correlation between OĀ VIĀ kinematics and disk rotation potentially implies that OĀ VIĀ alone does not serve as a useful tracer for galaxy gas accretion (Kacprzak etĀ al., 2019, 2025).
On the one hand, with the spatially extended distribution of OĀ VIĀ compared to low ions, it is perhaps unsurprising that the OĀ VIĀ kinematics do not seem to correlate with disk rotation as the low-ionization-state counterparts. In particular, because the CGM angular momentum remains roughly constant with radius, the outer CGM is expected to have a lower mean rotation velocity, resulting in a weaker correlation with the disk rotation. But on the other hand, in contrast to the large number of literatures that observationally confirmed the corotation between MgĀ IIĀ gas and the galaxy disk, only very few work has performed this kinematics comparison with OĀ VIĀ (Kacprzak etĀ al., 2019; Nateghi etĀ al., 2024; Kacprzak etĀ al., 2025).
In this paper, we build a sample of star-forming galaxies with quasar sightlines near the galaxy major axes and exclude galaxies with face-on disks. With quasar spectroscopy from the Cosmic Origin Spectrograph (COS) on the Hubble Space Telescope (HST) and galaxy rotation curve measurements, we analyze and compare the kinematics of the CGM traced by OĀ VIĀ and low ions (e.g., SiĀ II, SiĀ III) with the galaxy disk rotation. We describe the sample selection, data reduction, and measurements in SectionĀ 2. In SectionĀ 3, we present the results of the kinematics comparison among OĀ VIĀ gas, low-ionization-state gas, and disk rotation. We interpret our results and discuss their implications in SectionĀ 4. Finally, SectionĀ 5 summarizes the conclusions. Throughout the paper, we adopt the cosmology from Planck Collaboration etĀ al. (2016), with , , , and .
2 Data and Observations
We present the analyses of 18 galaxyāquasar pairs to study the kinematics of OĀ VIĀ and low-ionization-state absorbers and their relationship with galaxy disk rotation. We conducted quasar spectroscopy with HST/COS, including new spectra from our HST program and archival observations. We also acquired galaxy spectra from multiple ground-based telescopes. We describe our sample selection in SectionĀ 2.1. SectionĀ 2.2 discusses the galaxy spectroscopic observations and rotation curve measurements, and SectionĀ 2.3 describes the galaxy properties. In SectionĀ 2.4, we describe the HST/COS observations and measurements of absorption-line systems.
2.1 Sample of GalaxyāQuasar Pairs
Our study focuses star-forming galaxies with quasar sightlines near the major axes of the galaxies. The galaxyāquasar pairs were selected from published literature with known galaxy disk orientation; the disk geometry was obtained either from published work or from the Sloan Digital Sky Survey (SDSS; Blanton etĀ al., 2017). We restricted our sample to galaxies with inclined disks of inclination angles of ā, and each sightline lies within 45āĀ from the galaxy major axis, i.e., with azimuthal angle ā.
The compiled sample consists of 18 galaxyāquasar pairs with a median galaxy redshift of . TablesĀ 1 and 2 specify the coordinates and properties of galaxies, respectively, and TableĀ 3 lists the coordinates of the paired quasars. FigureĀ 1 shows the location of the quasar sightlines relative to the projected major and minor axes of the galaxies on the sky. The impact parameter of the quasar sightlines span between 20 and 280 kpc. The majority of the sightlines intersect the galaxies within 100 kpc, which corresponds to less than half of the virial radius () as indicated by the colors of the markers (see SectionĀ 2.3).
2.2 Galaxy Spectroscopy
We used the galaxy spectra to measure the systemic redshifts of the target galaxies and their rotation curves from the optical emission lines. We acquired new long-slit galaxy spectra with the Blue and Red Channel Spectrographs at the MMT Observatory. We also compiled previous observations (Ho etĀ al., 2017; Martin etĀ al., 2019; Kacprzak etĀ al., 2019) from the Low Resolution Imaging Spectrometer (LRIS; Oke etĀ al., 1995; Rockosi etĀ al., 2010) and the Echellette Spectrograph and Imager (ESI; Sheinis etĀ al., 2002) at the W.Ā M.Ā Keck Observatory and the Double Imaging Spectrograph (DIS) at the Apache Point Observatory (APO) 3.5 m telescope.222 Instrument specifications can be found in the manual written by Robert Lupton, which is available at http://www.apo.nmsu.edu/35m_operations/35m_manual/Instruments/DIS/DIS_usage.html#Lupton_Manual
We summarize the galaxy observations in TableĀ 1. SectionĀ 2.2.1 describes the new MMT observations, and SectionsĀ 2.2.2 and 2.2.3 briefly discuss the previous Keck and APO observations, respectively. SectionĀ 2.2.4 describes the measurements of the galaxy rotation curves from the emission lines (typically ) detected in the galaxy spectra.
Galaxy | Galaxy RA | Galaxy DEC | Instrument | Observing Date | Exposure Time | Reference |
(J2000) | (J2000) | (s) | ||||
J040150054047 | 04:01:50.48 | 05:40:47.07 | MMT/Red Channel | 2021 Jan 9 | 1800 | This work |
J080702+360141 | 08:07:02.27 | +36:01:41.15 | MMT/Blue Channel | 2020 Feb 28 | 2700 | This work |
J084356+261855 | 08:43:56.12 | +26:18:55.28 | MMT/Blue Channel | 2020 Feb 28 | 1800 | This work |
J085215+171137 | 08:52:15.36 | +17:11:37.07 | Keck/LRIS | 2014 May 2 | 2640 | Ho etĀ al. (2017) |
J091954+291345 | 09:19:54.11 | +29:13:45.34 | Keck/LRIS | 2015 Mar 22 | 1760 | Ho etĀ al. (2017) |
J094330+053118 | 09:43:30.67 | +05:31:18.11 | Keck/ESI | 2016 Jan 15 | 4500 | Kacprzak etĀ al. (2019)ā |
J103643+565119 | 10:36:43.44 | +56:51:19.00 | Keck/LRIS | 2015 Mar 22 | 1760 | Ho etĀ al. (2017) |
J104117+061018 | 10:41:17.80 | +06:10:18.88 | Keck/ESI | 2014 Apr 25 | 3300 | Kacprzak etĀ al. (2019)ā |
J113911135108 | 11:39:11.52 | 13:51:08.69 | Keck/ESI | 2016 Jan 15 | 1650 | Kacprzak etĀ al. (2019)ā |
J113909135053 | 11:39:09.80 | 13:50:53.08 | Keck/ESI | 2016 Jun 6 | 1800 | Kacprzak etĀ al. (2019)ā |
J123338+475757 | 12:33:38.87 | +47:57:57.62 | MMT/Red Channel | 2021 Jan 9 | 2400 | This work |
J132222+464546 | 13:22:22.46 | +46:45:46.18 | Keck/ESI | 2016 Jun 6 | 1500 | Kacprzak etĀ al. (2019)ā |
J135521+303320 | 13:55:21.20 | +30:33:20.41 | Keck/LRIS | 2014 May 3 | 3520 | Ho etĀ al. (2017) |
J142459+382113 | 14:24:59.82 | +38:21:13.37 | APO/DIS | 2016 Apr 2 | 5100 | Ho etĀ al. (2017) |
J154741+343350 | 15:47:41.46 | +34:33:50.86 | Keck/LRIS | 2014 May 3 | 1760 | Ho etĀ al. (2017) |
J155505+362848 | 15:55:05.27 | +36:28:48.40 | APO/DIS | 2016 Jul 4 | 3200 | Martin etĀ al. (2019) |
J160951+353838 | 16:09:51.62 | +35:38:38.58 | Keck/LRIS | 2014 May 2 | 1760 | Ho etĀ al. (2017) |
J225400+160925 | 22:54:00.58 | +16:09:25.83 | Keck/ESI | 2016 Jun 6 | 1200 | Kacprzak etĀ al. (2019)ā |
2.2.1 MMT Blue and Red Channel Observations
We conducted new galaxy spectroscopic observations using the MMT Blue and Red Channel Spectrographs. On February 28, 2020, we observed the galaxies with the Blue Channel and configured it with the 832 lines mm-1Ā grating blazed at 7790 Ć . On January 9, 2021, we used the Red Channel with the 1200 lines mm-1Ā grating blazed at 7700 Ć . For both observations, we applied the blue blocking filter R-63 to block the higher order light. We used the default detector binning, which provided a pixel scale of 03 for both channels and a wavelength dispersion of 0.72 Ć Ā pixel-1Ā and 0.80 Ć Ā pixel-1Ā for the Blue and Red Channel, respectively. We observed the target galaxies by aligning the 10-wide long-slit with the galaxy major axes. For both channels, the 10-wide slitlet provided a spectral resolution of 75ā85 km s-1Ā full width at half maximum intensity (FWHM).
The data were reduced using IRAF (Tody, 1986, 1993) following standard procedures. For individual exposures, the zero-point bias level was corrected for, and the fixed pattern noise was removed using exposures of the internal quartz lamp. Cosmic rays were removed using L.A. Cosmic (van Dokkum, 2001). Then, the frames were combined using a sigma-clipping algorithm to remove any remaining cosmic rays. A two-dimensional variance spectrum was computed when the exposures were combined. Frames of arc lamp exposures were used to calibrate the wavelength of the spectra, and the root-mean-square (rms) error of the dispersion solution was less than 15 km s-1. The wavelength solution was checked using the calibrated night-sky emission-line spectrum from the Ultraviolet and Visual Echelle Spectrograph (Hanuschik, 2003). Zero-point corrections were applied when necessary, which was caused by the difference in rotator angles between the science exposures and the calibration frames. This was followed by a heliocentric correction to correct for the earthās seasonal motion.
From the two-dimensional galaxy spectra, we measured the galaxy rotation curves from the emission lines (SectionĀ 2.2.4). We also extracted an one-dimensional spectrum for each galaxy and determined the galaxy systemic redshift from the redshifted wavelengths of the , [NĀ II] 6548, 6583, and [SĀ II] 6716, 6731Ā emission lines.
2.2.2 Keck Observations
The Keck/LRIS observations were described in detail in Ho etĀ al. (2017) and Martin etĀ al. (2019). The LRIS double spectrograph was configured using the D500 dichroic, 1200 lines mm-1Ā grism blazed at 3400Ć Ā (LRISb) and the 900 lines mm-1Ā grism blazed at 5500Ć Ā (LRISr). The galaxy emission lines for rotation curve measurements were captured by LRISr. The detectors were binned by two in both spatial and dispersion directions, which produced a spatial scale of and a dispersion of 1.06 Ć Ā pixel-1. The 10-wide long-slit was either aligned with the position angle of the major axis of the target galaxies, or masks of 10-wide slitlets were designed to include nearby galaxies in the field. The 10-wide slitlets provided an average resolution of 75ā105 km s-1Ā FWHM. The data reduction procedures for the LRISr spectra were explained in Ho etĀ al. (2017) and were similar to those for reducing the MMT spectra described in SectionĀ 2.2.1.
The Keck/ESI observations were described in detail in Kacprzak etĀ al. (2019). In brief, the galaxies were observed using the 10-wide, 20ā²ā²-long ESI slit, and the slitlet was aligned with the major axes of individual galaxies. The detector was binned by two in both the spatial and spectral directions. This produced a dispersion of 22 km s-1Ā pixel-1 and pixel scales of 027 to 034 over the echelle orders that covered the galaxy optical emission lines. The resulting spectral resolution from the 10-wide slitlet was 90 km s-1Ā FWHM. Following the data reduction procedures as outlined in Kacprzak etĀ al. (2019), we re-reduced the ESI spectra and re-examined the galaxy emission lines. The new two-dimensional spectra were adopted to measure the galaxy rotation curves (SectionĀ 2.2.4).
2.2.3 APO DIS Observations
For galaxies J142459+38211 and J155505+362848, Ho etĀ al. (2017) and Martin etĀ al. (2019) conducted long-slit spectroscopy with APO/DIS to measure the galaxy rotation curves from the Ā emission line. The 15-wide long-slit was aligned with the major axis of each galaxy. The blue and red channels of DIS were configured with the 1200 lines mm-1Ā blazed at 4400Ć Ā and 7300Ć , respectively. The Ā emission was detected by the red spectra, which had a spatial scale and dispersion of 040 and 0.58 Ć Ā pixel-1, respectively. The DIS spectra were reduced as discussed in Ho etĀ al. (2017) and followed the same procedures as we reduced the MMT spectra. The spectral resolution was measured using the arclampsā lines as 50 km s-1Ā FWHM.
2.2.4 Galaxy Rotation Curves
We generate the galaxy positionāvelocity (PV) maps by fitting a Gaussian profile to the galaxy emission lines at different spatial locations on the two-dimensional galaxy spectra. Following Ho etĀ al. (2017), we convert the PV maps to rotation curve measurements by deprojecting the position and velocity measured along the slit onto circular orbits on the disk plane. For a slitlet at an angle from the galaxy major axis, we calculate the galactocentric radius from the projected distance along the slit using
where represents the disk inclination angle. We convert the line-of-sight velocity to rotation velocity on the disk by
In the middle column of FigureĀ 2, we show the line-of-sight velocity and the rotation velocity of individual galaxies, which are represented by the cyan solid squares and black open circles, respectively.
To model the intrinsic rotation curves of the galaxies, we define a grid of rotation curves over a range of asymptotic rotation velocities and turnover radii using an arctangent model,
For each galaxy, we project the models onto the slit, convolve these projected models with a Gaussian model for the seeing disk, and compare them to the PV maps measured from the galaxy emission lines. We minimize the fit residual to estimate Ā and for individual galaxies. In the middle column of FigureĀ 2, the seeing-convolved model for each galaxy before and after the projection are shown using the black dashed and cyan solid curves, respectively. TableĀ 2 lists the adopted Ā and for individual galaxies.
2.3 Galaxy Properties
We summarize the properties of the selected galaxies in TableĀ 2. The systemic redshifts of the galaxies are spectroscopically determined from the optical emission lines either from the new MMT spectroscopy or from previous work, and we specify the references in the Table. SectionsĀ 2.3.1 and 2.3.2 describe the galaxy disk geometry and the stellar and halo masses, respectively. Then, we discuss the galaxy environments in SectionĀ 2.3.3.
Quasar | Galaxy | ||||||||||
() | () | (kpc) | (km s-1) | (kpc) | (ā) | (ā) | (kpc) | ||||
J040148054056 | J040150054047 | 0.21966 | 9.9aaAdopted from Werk etĀ al. (2012). Galaxy stellar masses are rescaled from Salpeter to Chabrier IMF. | 11.6 | 168 | 140 | 4.8 | 58bbDerived from SDSS DR17 PhotoObj catalog for -band images. Inclination angle of the galactic disk is calculated from the disk axis ratio using the Hubble formula (Hubble, 1926) with . Azimuthal angle of the quasar sightline is calculated from the galaxy major-axis position angle in the catalog. | 24bbDerived from SDSS DR17 PhotoObj catalog for -band images. Inclination angle of the galactic disk is calculated from the disk axis ratio using the Hubble formula (Hubble, 1926) with . Azimuthal angle of the quasar sightline is calculated from the galaxy major-axis position angle in the catalog. | 89 | 0.53 |
J080704+360353 | J080702+360141 | 0.088061 | 10.9ccAdopted from Heckman etĀ al. (2017). Galaxy stellar masses are rescaled from Kroupa to Chabrier IMF. | 12.7 | 426 | 250 | 1.4 | 71bbDerived from SDSS DR17 PhotoObj catalog for -band images. Inclination angle of the galactic disk is calculated from the disk axis ratio using the Hubble formula (Hubble, 1926) with . Azimuthal angle of the quasar sightline is calculated from the galaxy major-axis position angle in the catalog. | 32bbDerived from SDSS DR17 PhotoObj catalog for -band images. Inclination angle of the galactic disk is calculated from the disk axis ratio using the Hubble formula (Hubble, 1926) with . Azimuthal angle of the quasar sightline is calculated from the galaxy major-axis position angle in the catalog. | 231 | 0.54 |
J084349+261910 | J084356+261855 | 0.11284 | 10.5ccAdopted from Heckman etĀ al. (2017). Galaxy stellar masses are rescaled from Kroupa to Chabrier IMF. | 12.1 | 261 | 190 | 1.5 | 55bbDerived from SDSS DR17 PhotoObj catalog for -band images. Inclination angle of the galactic disk is calculated from the disk axis ratio using the Hubble formula (Hubble, 1926) with . Azimuthal angle of the quasar sightline is calculated from the galaxy major-axis position angle in the catalog. | 42bbDerived from SDSS DR17 PhotoObj catalog for -band images. Inclination angle of the galactic disk is calculated from the disk axis ratio using the Hubble formula (Hubble, 1926) with . Azimuthal angle of the quasar sightline is calculated from the galaxy major-axis position angle in the catalog. | 184 | 0.71 |
J085215+171143 | J085215+171137 | 0.16921ddAdopted from Martin etĀ al. (2019). | 9.8ddAdopted from Martin etĀ al. (2019). | 11.5 | 162 | 200 | 2.7 | 77ddAdopted from Martin etĀ al. (2019). | 23ddAdopted from Martin etĀ al. (2019). | 20 | 0.12 |
J091954+291408 | J091954+291345 | 0.23288ddAdopted from Martin etĀ al. (2019). | 10.5ddAdopted from Martin etĀ al. (2019). | 12.2 | 264 | 250 | 3.4 | 73ddAdopted from Martin etĀ al. (2019). | 15ddAdopted from Martin etĀ al. (2019). | 88 | 0.34 |
J094331+053131 | J094330+053118 | 0.353052eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 9.4aaAdopted from Werk etĀ al. (2012). Galaxy stellar masses are rescaled from Salpeter to Chabrier IMF. | 10.3 | 128 | 105 | 1.5 | 44eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 8eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 99 | 0.77 |
J103640+565125 | J103643+565119 | 0.13629ddAdopted from Martin etĀ al. (2019). | 9.9ddAdopted from Martin etĀ al. (2019). | 11.6 | 174 | 210 | 1.7 | 58ddAdopted from Martin etĀ al. (2019). | 21ddAdopted from Martin etĀ al. (2019). | 58 | 0.33 |
J104116+061016 | J104117+061018 | 0.442173eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 10.3 | 12.0eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 203 | 320 | 1.2 | 50eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 4eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 57 | 0.28 |
J113910135043 | J113911135108 | 0.204194eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 10.1 | 11.7eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 184 | 175 | 2.4 | 82eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 6eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 96 | 0.52 |
J113910135043 | J113909135053 | 0.319255eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 10.2 | 11.9eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 197 | 205 | 2.9 | 83eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 39eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 77 | 0.39 |
J123335+475800 | J123338+475757 | 0.22171 | 10.6aaAdopted from Werk etĀ al. (2012). Galaxy stellar masses are rescaled from Salpeter to Chabrier IMF. | 12.2 | 267 | 225 | 1.5 | 74bbDerived from SDSS DR17 PhotoObj catalog for -band images. Inclination angle of the galactic disk is calculated from the disk axis ratio using the Hubble formula (Hubble, 1926) with . Azimuthal angle of the quasar sightline is calculated from the galaxy major-axis position angle in the catalog. | 5bbDerived from SDSS DR17 PhotoObj catalog for -band images. Inclination angle of the galactic disk is calculated from the disk axis ratio using the Hubble formula (Hubble, 1926) with . Azimuthal angle of the quasar sightline is calculated from the galaxy major-axis position angle in the catalog. | 141 | 0.53 |
J132222+464546 | J132222+464546 | 0.214431eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 10.6aaAdopted from Werk etĀ al. (2012). Galaxy stellar masses are rescaled from Salpeter to Chabrier IMF. | 12.2 | 277 | 180 | 1.8 | 58eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 14eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 40 | 0.15 |
J135522+303324 | J135521+303320 | 0.20690ddAdopted from Martin etĀ al. (2019). | 9.8ddAdopted from Martin etĀ al. (2019). | 11.5 | 157 | 155 | 2.8 | 72ddAdopted from Martin etĀ al. (2019). | 2ddAdopted from Martin etĀ al. (2019). | 78 | 0.50 |
J142501+382100 | J142459+382113 | 0.21295ddAdopted from Martin etĀ al. (2019). | 10.2ddAdopted from Martin etĀ al. (2019). | 11.7 | 190 | 190 | 2.2 | 61ddAdopted from Martin etĀ al. (2019). | 7ddAdopted from Martin etĀ al. (2019). | 83 | 0.44 |
J154741+343357 | J154741+343350 | 0.18392ddAdopted from Martin etĀ al. (2019). | 10.0ddAdopted from Martin etĀ al. (2019). | 11.6 | 173 | 175 | 1.6 | 59ddAdopted from Martin etĀ al. (2019). | 4ddAdopted from Martin etĀ al. (2019). | 26 | 0.15 |
J155504+362847 | J155505+362848 | 0.18926ddAdopted from Martin etĀ al. (2019). | 10.1ddAdopted from Martin etĀ al. (2019). | 11.7 | 188 | 255 | 3.6 | 55ddAdopted from Martin etĀ al. (2019). | 45ddAdopted from Martin etĀ al. (2019). | 35 | 0.19 |
J160951+353843 | J160951+353838 | 0.28940ddAdopted from Martin etĀ al. (2019). | 9.9ddAdopted from Martin etĀ al. (2019). | 11.6 | 164 | 45 | 2.7 | 65ddAdopted from Martin etĀ al. (2019). | 25ddAdopted from Martin etĀ al. (2019). | 25 | 0.15 |
J225357+160853 | J225400+160925 | 0.390013eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 10.5 | 12.2eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 238 | 175 | 3.3 | 76eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 24eeAdopted from Kacprzak etĀ al. (2019). With their halo masses, the stellar masses are re-derived using the stellar massāhalo mass relation from Behroozi etĀ al. (2010). | 284 | 1.19 |
Note. ā (1) Name of the quasar. (2) Name of the galaxy. (3) Galaxy systemic redshift. (4) Galaxy stellar mass, rescaled to Chabrier IMF when necessary following Madau & Dickinson (2014). (5) Halo mass, derived using the stellar massāhalo mass relation in Behroozi etĀ al. (2010) except for those adopted from Kacprzak etĀ al. (2019). (6) Halo virial radius, defined by the overdensity in Bryan & Norman (1998) with respect to the critical density evaluated at the galaxy systemic redshift. (7) Asymptotic rotation speed of the galaxy disk. (8) Turnover radius from the arctangent rotation curve model. (9) Inclination angle of galactic disk. (10) Azimuthal angle, the angle between the quasar sightline and the galaxy major axis. (11) Sightline impact parameter, the projected distance between the centers of the galaxy and quasar. (12) Sightline impact parameter normalized by the halo virial radius.
2.3.1 Galaxy Disk Geometry
We adopted measurements of galaxy disk geometry either from previous work or from the SDSS DR17 catalog (Blanton etĀ al., 2017). TableĀ 2 lists the adopted values and the references for the disk inclination angles and the azimuthal angles of the quasar sightlines with respect to the galaxy major axes. Both Martin etĀ al. (2019) and Kacprzak etĀ al. (2019) measured the galaxy disk geometry using high-resolution images: the former obtained -band images from the Keck/NIRC2 camera with Laser Guide Star Adaptive Optics system (Wizinowich etĀ al., 2006), and the latter collected archival HST images from the Advanced Camera for Surveys (ACS) and Wide Field and Planetary Camera 2 (WFPC2). For the remaining galaxies, we used the -band photometry in the SDSS PhotoObj catalog to derive the azimuthal angle and the inclination angle from the axis ratio333 We applied the Hubble formula (Hubble, 1926) with . . We show the images of individual galaxies with the same instruments and bandpasses used for measuring the parameters of the disk geometry in FigureĀ 2.
2.3.2 Stellar Masses, Halo Masses, and Virial Radii
The stellar masses Ā of individual galaxies were obtained from previous work and were rescaled to the Chabrier initial mass function (IMF; Chabrier, 2003) following Madau & Dickinson (2014). Stellar masses extracted from Martin etĀ al. (2019) were in Chabrier IMF, whereas Werk etĀ al. (2012) assumed the Salpeter IMF (Salpeter, 1955), and Heckman etĀ al. (2017) adopted the Kroupa IMF (Kroupa, 2001) as in the MPA-JHU Value-Added Galaxy Catalog. Then, the halo mass Ā of each galaxy was calculated using the stellar massāhalo mass (SM-HM) relation derived from abundance matching in Behroozi etĀ al. (2010). The virial radius Ā was obtained following Ho etĀ al. (2017) using the overdensity defined by Bryan & Norman (1998).
The above procedures were applied for 14 of the 18 galaxies, for which the stellar masses were derived from multi-band galaxy photometry using the spectral energy distribution fitting technique. But for four galaxies from Kacprzak etĀ al. (2019), instead of first obtaining the stellar masses and then calculate the halo masses, they derived the halo masses from the -band Vega magnitudes of the galaxies following the method in Churchill etĀ al. (2013) (also see Ng etĀ al., 2019). We used their halo masses and calculated the stellar masses of the galaxies using the SM-HM relation from Behroozi etĀ al. (2010). Then, we derived the virial radii from the halo masses following the same procedures as we did for other galaxies.
TableĀ 2 lists the stellar mass, halo mass, and virial radius of individual galaxies, and we label the references of the masses. The median Ā and Ā are 10.14 and 11.72, respectively. The Table also shows the sightline impact parameter normalized by the halo virial radius . All except one sightline intersect the galaxies within the virial radius (), and most of the sightlines lie within 0.5Ā from the galaxies (also see FigureĀ 1).
2.3.3 Galaxy Environments and Potential Interactions
The group environment of galaxies affect their CGM properties, such as the covering fraction and velocity dispersion, and the CGM becomes kinematically complex (Johnson etĀ al., 2015; Pointon etĀ al., 2017; Nielsen etĀ al., 2018; Dutta etĀ al., 2025). It also leads to uncertainties in identifying the host galaxies associated with the absorption detected in the sightlines. Because we generally lack spectroscopic redshifts for most galaxies in the our fields, we cannot classify the group environment as detailed as in Johnson etĀ al. (2013) and Chen etĀ al. (2020). Nevertheless, using SDSS and SIMBAD (Wenger etĀ al., 2000), we searched for galaxies near our targets to confirm there were no other galaxies with spectroscopic or photometric redshifts comparable to our target galaxies, especially for galaxies with similar -band magnitudes and/or impact parameters as our targets.
However, the only exception is the target galaxy J091954+291345 at . We found that two red galaxies identified by the SHELS galaxy redshift survey (Geller etĀ al., 2014) have spectroscopic redshifts similar to our target. At the northeast of our target galaxy, a bright red galaxy 44 away (17 kpc) has a redshift of 0.23208, i.e., km s-1Ā relative to our galaxy. The other fainter, red galaxy has a redshift of 0.23179, i.e., km s-1, at 87 (33 kpc) south of our target. Hence, these three galaxies form a group with velocity differences within a few hundred km s-1. While we present the spectroscopic measurements for this target galaxy and the absorption lines measured in the quasar sightline, we exclude this system from the CGM kinematics analyses in later sections.
For galaxy J160951+353838, it shows a tail-like structure towards the northeast (also see Ho etĀ al. 2017). This tail can be seen towards the upper right in the galaxy image stamp in FigureĀ 2. The galaxy rotation curve also shows irregularity in the same direction. The presence of the tail and the galaxy kinematics together suggest interaction with another galaxy. However, we do not find any bright galaxies with comparable photometric redshift that could have interacted with our target galaxy within its virial radius. While the presence of a less massive galaxy or satellite potentially explains the irregular galaxy morphology and kinematics, because our target is not in a close group as J091954+291345, we do not exclude this galaxy in the CGM kinematics analysis.
2.4 Quasar Spectroscopy with HST/COS
The far-UV (FUV) spectroscopy of the background quasars of individual galaxyāquasar pairs were either from our HST/COS program (ID: 15866, PI: Ho) or from archival COS data. These COS spectra were obtained using the medium-resolution grating, G130M and/or G160M, with a spectral resolution of . TableĀ 3 lists the details of the quasar observations with COS.
For each quasar, the raw data from COS were processed by the CALCOS pipeline (v3.3.9), and the one-dimensional spectra (x1d files) were retrieved from MAST444http://archive.stsci.edu. Custom software was used to align the one-dimensional spectra using the Milky Way absorption lines and coadd the aligned spectra. Because COS FUV spectra were oversampled, the coadded spectra were binned by three pixels to increase the signal-to-noise ratio. Finally, each coadded spectrum was continuum-normalized by fitting the quasar continuum with an Akime Spline at the absorption-free regions using the linetools555https://github.com/linetools/linetools python package.
Quasar | Quasar RA | Quasar DEC | Grating | Exposure Time | Program ID | |
(J2000) | (J2000) | (s) | ||||
J040148054056 | 04:01:48.98 | 05:40:56.5 | 0.570 | G130M/G160M | 5377/5912 | 11598 |
J080704+360353 | 08:07:04.88 | +36:03:53.5 | 0.468 | G130M/G160M | 5378/8612 | 13862 |
J084349+261910 | 08:43:49.75 | +26:19:10.7 | 0.257 | G130M | 2194 | 13862 |
J085215+171143 | 08:52:15.34 | +17:11:43.8 | 0.562 | G130M | 10810 | 15866 |
J091954+291408 | 09:19:54.28 | +29:14:08.3 | 0.722 | G130M | 4933 | 15866 |
J094331+053131 | 09:43:31.61 | +05:31:31.4 | 0.564 | G130M/G160M | 3662/3945 | 11598 |
J103640+565125 | 10:36:40.74 | +56:51:25.9 | 1.264 | G130M | 8641 | 15866 |
J104116+061016 | 10:41:17.16 | +06:10:16.8 | 1.270 | G160M | 14183 | 12252 |
J113910135043 | 11:39:10.70 | 13:50:43.6 | 0.557 | G130M | 7669 | 12275 |
J123335+475800 | 12:33:35.07 | +47:58:00.4 | 0.382 | G130M/G160M | 3885/4218 | 11598 |
G130M | 2044 | 13033 | ||||
J132222+464546 | 13:22:22.67 | +46:45:35.2 | 0.374 | G130M/G160M | 3902/4249 | 11598 |
G130M | 2044 | 13033 | ||||
J135522+303324 | 13:55:22.89 | +30:33:24.7 | 0.382 | G130M | 8092 | 15866 |
J142501+382100 | 14:25:01.46 | +38:21:00.5 | 1.145 | G130M | 8122 | 15866 |
J154741+343357 | 15:47:41.88 | +34:33:57.3 | 0.928 | G130M | 5197 | 15866 |
J155504+362847 | 15:55:04.39 | +36:28:47.9 | 0.714 | G130M/G160M | 5151/6030 | 11598 |
J160951+353843 | 16:09:51.81 | +35:38:43.7 | 0.828 | G130M | 10639 | 15866 |
J225357+160853 | 22:53:57.74 | +16:08:53.5 | 0.859 | G130M/G160M | 4473/4514 | 13398 |
2.5 Absorption-line Measurements: Voigt Profile Fitting and Column-density-weighted Velocity
We identified absorption-line systems within km s-1Ā of the systemic velocity of the target galaxies. We modeled individual velocity components with Voigt profiles and adopted the atomic data from Morton (2003). The modeled Voigt profile characterized the velocity centroid , the line width, i.e., the Doppler parameter , and the column density of individual velocity component. Using the VoigtFit python package (Krogager, 2018), we convolved the models with the COS line-spread function corresponding to the COS lifetime position when the observation was conducted. We compared the convolved line profiles to the data, and different transitions of the same ionic species were fitted jointly. For each ionic species, we created different sets of models with different number of velocity components, and the best-fit solution of each model set was determined using a minimization algorithm. Finally, we adopted the best-fit model that required the minimum number of velocity components to produce a reasonable fit and with the reduced- value closed to one.
The right column of FigureĀ 2 shows the OĀ VIĀ absorption lines relative to the systemic redshifts of individual target galaxies. The fitted Voigt profiles of individual velocity components and the overall OĀ VIĀ absorption are shown in blue dashed and solid curves, respectively, and the blue ticks (top) mark the velocity centroids of individual components. TableĀ 4 lists the parameters of the Voigt profiles of the fitted OĀ VIĀ absorption.666 The Figures of the fitted line profiles of other ions and Voigt profile parameters are shown in FigureĀ 13 and TableĀ 6, respectively, in the Appendix. Only the J103640565125 sightline did not detect OĀ VIĀ around the target galaxy. In that case, we list the column density upper limit; this was derived from the linear section of the curve-of-growth and the upper limit of the rest-frame equivalent width of OĀ VIĀ 1031, measured using a velocity window that was three times the resolution element centered at the galaxy systemic velocity.
Quasar | Galaxy | |||||
(km s-1) | (km s-1) | () | (km s-1) | |||
J040148054056 | J040150054047 | 0.21966 | ||||
J080704+360353 | J080702+360141 | 0.088061 | ||||
J084349+261910 | J084356+261855 | 0.11284 | ||||
J085215+171143 | J085215+171137 | 0.16921 | ||||
J094331+053131 | J094330+053118 | 0.353052 | ||||
J103640+565125 | J103643+565119 | 0.13629 | ⦠| ⦠| ā | ⦠|
J104116+061016 | J104117+061018 | 0.442173 | ||||
J113910135043 | J113911135108 | 0.204194 | ||||
J113910135043 | J113909135053 | 0.319255 | ||||
J123335+475800 | J123338+475757 | 0.22171 | ||||
J132222+464546 | J132222+464546 | 0.214431 | ||||
J135522+303324 | J135521+303320 | 0.20690 | ||||
J142501+382100 | J142459+382113 | 0.21295 | ||||
J154741+343357 | J154741+343350 | 0.18392 | ||||
J155504+362847 | J155505+362848 | 0.18926 | ||||
J160951+353843 | J160951+353838 | 0.28940 | ||||
J225357+160853 | J225400+160925 | 0.390013 | ||||
Excluded from OĀ VIĀ kinematics analysis due to group environment | ||||||
J091954+291408 | J091954+291345 | 0.23288 | ||||
Note. ā (1) Name of the quasar. (2) Name of the galaxy. (3) Galaxy systemic redshift. (4) Velocity centroid of individual component. (5) Line width, i.e., Doppler parameter, of the velocity component. (6) Logarithm of the column density of the velocity component. (7) Column-density-weighted velocity.
For each sightline,
we calculate the column-density-weighted velocity
of the overall absorption.
Using the results from Voigt profile fitting,
we perturb the fitted velocity centroid and
column density of each velocity component by
adding a random value within their 1- error bars,
and we repeat this process 1000 times.
During the -th iteration,
we determine the weighted velocity by
where indicates the -th velocity component, and and represent the perturbed velocity centroid and column density, respectively. From the resultant distribution of the weighted velocity, we adopt the median as the column-density-weighted velocity and the 68% confidence interval as the uncertainties. We show the OĀ VIĀ column-density-weighted velocity Ā for each sightline in the rightmost column of TableĀ 4.
3 CGM Kinematics and Galaxy Disk Rotation
With our sample of quasar sightlines near the major axes of galaxies, we examine the relationship between CGM kinematics and galaxy disk rotation. SectionĀ 3.1 and SectionĀ 3.2 focus on the OĀ VIĀ CGM and the low-ionization-state gas, respectively. In SectionĀ 3.3, we then compare the kinematics of the low ions and OĀ VIĀ detected in individual sightlines. In our analyses, we exclude galaxy J091954+291345 due to its close group environment (SectionĀ 2.3.3) and focus on the remaining 17 galaxyāquasar pairs.
3.1 Kinematics of OĀ VIĀ Gas
Among the 17 galaxyāquasar pairs, all sightlines except J103640+565125 have detected OĀ VIĀ around the host galaxies. FigureĀ 5 compares the kinematics of the OĀ VIĀ gas and that of the galaxy disk rotation. The circles and the dashed curves represent the velocity centroids of individual OĀ VIĀ velocity components and the galaxy rotation curves, respectively. With similar numbers of OĀ VIĀ velocity components that corotate (21) and counter-rotate (15) with the disk (TableĀ 5), it seems to suggest a lack of correlation between disk rotation and the kinematics of individual OĀ VIĀ velocity components. However, except for the sightline at the highest impact parameter of 284 kpc, the column-density-weighted velocities (black crosses) suggest that the bulk of OĀ VIĀ gas either corotates with the disk or is consistent with being at the systemic velocity of the host galaxies (within the grey band).777 Velocities smaller than the measurement uncertainties of the galaxy redshift of 25 km s-1Ā are considered as being consistent with the galaxy systemic velocity. In other words, we rarely found sightlines with net counter-rotating OĀ VIĀ absorbers.
In the bottom panel of FigureĀ 5, we normalized the velocities and impact parameters by the asymptotic rotation speeds and virial radii of the galaxies, respectively. The velocity centroids of most OĀ VIĀ components (34 out of 41) are smaller than the asymptotic rotation speed, i.e., . Sightlines with OĀ VIĀ components exceeding Ā typically have large impact parameters of 0.5. The only exception is the sightline around host galaxy J160951+353838, but this galaxy has a tail-like structure and an abnormally low asymptotic rotation speed of 45 km s-1Ā (SectionĀ 2.3.3). This explains the high values even though the measured OĀ VIĀ velocities in km s-1Ā are comparable to those in other sightlines.
OĀ VIĀ Category | Kinematics Class | Ratio | |
All Sightlines (exclude J091929) | |||
All | Any | 41 | ⦠|
Corotate | 21 | 0.51 | |
Systemic Velocity | 5 | 0.12 | |
Counter-rotate | 15 | 0.37 | |
Sightlines with SiĀ IIĀ and SiĀ IIIĀ Spectral Coverage ()ā | |||
All | Any | 29 | ⦠|
Corotate | 16 | 0.55 | |
Systemic Velocity | 3 | 0.10 | |
Counter-rotate | 10 | 0.34 | |
With Si II match | Any | 10 | ⦠|
Corotate | 9 | 0.90 | |
Systemic Velocity | 1 | 0.10 | |
Counter-rotate | 0 | 0 | |
No Si II match | Any | 19 | ⦠|
Corotate | 7 | 0.37 | |
Systemic Velocity | 2 | 0.11 | |
Counter-rotate | 10 | 0.53 | |
With Si III match | Any | 14 | ⦠|
Corotate | 10 | 0.71 | |
Systemic Velocity | 2 | 0.14 | |
Counter-rotate | 2 | 0.14 | |
No Si III match | Any | 15 | ⦠|
Corotate | 6 | 0.40 | |
Systemic Velocity | 1 | 0.07 | |
Counter-rotate | 8 | 0.53 |
Note. ā (1) Category of OĀ VIĀ velocity components. (2) Classification of OĀ VIĀ velocity components based on their Doppler sign relative to the disk rotation at the quasar side of the galaxy disk. (3) Number of OĀ VIĀ velocity components. (4) The ratio of the number of OĀ VIĀ velocity components between the kinematics class and the total within the same OĀ VIĀ category.
3.2 Corotation of the Low-ionization-state Gas with the Disk
Low-ionization-state ions trace cooler gas compared to OĀ VIĀ due to their lower ionization potentials (IP), e.g., SiĀ IIĀ and SiĀ IIIĀ with IP of 8.2 eV and 16.3 eV, respectively, compared to 113.9 eV for producing OĀ VI. For our sample, only 13 sightlines have COS spectra that cover the redshifted wavelengths of SiĀ II888 We focus on the SiĀ II 1190, 1193Ā and SiĀ II 1260Ā transitions, because other SiĀ IIĀ transition lines within the spectral coverage are either too weak to be detected or blended with other absorption lines. and SiĀ IIIĀ of the host galaxies, including the J103640+565125 sightline without OĀ VIĀ detection. Among these 13 sightlines, SiĀ IIIĀ exhibits a higher detection rate compared to SiĀ II, with 7 and 11 sightlines showing SiĀ IIĀ and SiĀ IIIĀ detection, respectively.
FigureĀ 6 compares the kinematics of the SiĀ IIĀ (left) and SiĀ IIIĀ (right) gas to disk rotation. Except for the sightline at 184 kpc from galaxy J084356+261855, six out of the seven SiĀ II-detected sightlines show net corotating SiĀ IIĀ absorption, and only two (out of 21) SiĀ IIĀ velocity components counter-rotate with the disk. Net counter-rotating SiĀ IIIĀ absorbers are also rare, but the fraction of SiĀ IIIĀ velocity components that counter-rotate with the disk becomes higher compared to that of SiĀ II.
FigureĀ 7 shows the fraction of non-counter-rotating velocity components of SiĀ II, SiĀ III, and OĀ VI, i.e., the gas either corotates with the disk or consistent with being at the systemic velocity of the host galaxy.999 Both black hatched and maroon filled histograms represent OĀ VI, but the former includes all sightlines, whereas the latter only includes the 13 sightlines with SiĀ IIĀ and SiĀ IIIĀ spectral coverage. Considering all sightlines regardless of their impact parameters (āAllā), SiĀ IIĀ has the highest fraction of non-counter-rotating velocity components of 90%. The fraction drops with increasing ionization potential, for which only 75% (60%) of the SiĀ IIIĀ (OĀ VI) velocity components are not counter-rotating. If we only focus on the sightlines with impact parameters below 100 kpc, the fraction of non-counter-rotating velocity components for all three ions become comparable (between 75% and 95%). However, in sightlines beyond 100 kpc from the host galaxies, OĀ VIĀ shows a significantly lower non-counter-rotating fraction compared to SiĀ IIĀ and SiĀ IIIĀ in the same impact parameter range and also compared to the OĀ VIĀ detected at small impact parameters. Therefore, in contrast to low-ionization-state gas that rarely counter-rotates with the disk, whether individual velocity components of the OĀ VIĀ gas counter-rotate depends on the sightline impact parameter, with sightlines at larger impact parameters more frequently detect counter-rotating OĀ VIĀ gas.
3.3 Comparison Between OĀ VIĀ and Low-ionization-state Gas Kinematics
SectionsĀ 3.1 and 3.2 have shown that the low-ionization-state gas traced by SiĀ IIĀ and SiĀ IIIĀ rarely counter-rotate with the disk compared to OĀ VI, suggesting that OĀ VIĀ gas exhibits different kinematics compared to the lower ionization counterparts. In FigureĀ 8, we compare the Voigt profile fits of OĀ VIĀ to those of SiĀ IIĀ and SiĀ IIIĀ measured in individual sightlines.101010 For the J154741343357 sightline, the SiĀ IIIĀ absorption-line profile is incomplete, because it falls at the reddest end of the spectrum. The J103640+565125 sightline is not shown due to OĀ VIĀ non-detection. We arrange the panels in increasing impact parameter of the quasar sightlines. First, SiĀ IIĀ and SiĀ IIIĀ are more frequently detected in sightlines at low impact parameters, and the two sightlines beyond 200 kpc from the host galaxies neither detect SiĀ IIĀ nor SiĀ III. Secondly, the line profiles of different ions have complex component structures, especially sightlines closer to the galaxies tend to have more velocity components. While some OĀ VIĀ components have velocities comparable to their low-ionization-state counterparts, some do not affiliate with the low ions. But regardless, the overall absorption of different ions typically occupy similar velocity ranges.
Motivated by FigureĀ 8 that shows that only some OĀ VIĀ components have velocities comparable to those of the low ions, we examine how the kinematics association between OĀ VIĀ and low ions (or the lackthereof) relates to the comparison of OĀ VIĀ kinematics and disk rotation. In each sightline, we attempt to match individual OĀ VIĀ velocity components with those of SiĀ IIĀ and SiĀ IIIĀ by minimizing the difference of their velocity centroids. We consider an OĀ VIĀ velocity component to be successfully matched with a SiĀ IIĀ or SiĀ IIIĀ component if the velocity difference is under 30 km s-1, a threshold that includes the uncertainties in defining the galaxy systemic velocity, the wavelength calibration of the COS spectra111111 According to the āCycle 26 COS/FUV Spectroscopic Sensitivity Monitor (ISR 2020-06)ā, the COS wavelength zeropoint uncertainties are 3 pixels, i.e., km s-1. , and the typical uncertainty of the velocity centroid from Voigt profile fitting.121212 While increasing (decreasing) the 30 km s-1Ā threshold value will increase (reduce) the number of low-ion matched OĀ VIĀ velocity components, it does not affect the main conclusions from the kinematics comparison between OĀ VIĀ components with and without low-ion matches. Each SiĀ IIĀ or SiĀ IIIĀ velocity component can only be matched with one OĀ VIĀ component (and vice versa). Hence, if more than one SiĀ IIĀ or SiĀ IIIĀ component satisfy the matching criteria with OĀ VI, then we select the one with the smallest velocity difference.
Analogous to FigureĀ 5, FigureĀ 9 compares the OĀ VIĀ kinematics with galaxy disk rotation by showing the 12 sightlines with OĀ VIĀ detection and spectral coverage of SiĀ IIĀ and SiĀ III. Each marker represents the velocity centroid of an OĀ VIĀ component, and the yellow diamonds (top) and cyan squares (bottom) highlight the SiĀ II- and SiĀ III-matched OĀ VIĀ components, respectively. Both panels clearly show that SiĀ II- or SiĀ III-matched OĀ VIĀ velocity components rarely counter-rotate with the galaxy disk (also see TableĀ 5). In fact, none of the SiĀ II-matched OĀ VIĀ velocity components counter-rotates (0/10), in contrast to the 10/19 (53%) no-SiĀ II-matched OĀ VIĀ components that counter-rotate with the disk. While counter-rotating SiĀ III-matched OĀ VIĀ components exist (2/14; 14%), that occurs less frequently compared to the OĀ VIĀ components without a SiĀ IIIĀ matched (8/15; 53%).
Not only the SiĀ II- and SiĀ III-matched OĀ VIĀ components rarely counter-rotate, but they also preferentially corotate with the disk instead of being at the galaxy systemic velocity. Specifically, 9/10 (90%) of the SiĀ II-matched OĀ VIĀ components corotate with the disk in contrast to only 7/19 (37%) of the OĀ VIĀ components without a SiĀ IIĀ match. The contrast becomes less prominent for SiĀ III. But still, a higher fraction (10/14; 71%) of the SiĀ III-matched OĀ VIĀ components corotate with the disk compared to that of the no-SiĀ III-matched OĀ VIĀ components (6/15; 40%). Our results show that the OĀ VIĀ velocity components with and without low-ion matches have different kinematics properties and possibly trace gas that resides at different 3D distances from the host galaxies. Because OĀ VIĀ velocity components with matched low-ion are more likely to corotate with the disk, a plausible scenario is that the gas resides on an extended disk plane. We further discuss the implications of our results in SectionĀ 4.
4 Discussion
In SectionĀ 3, we have presented the kinematics of the OĀ VIĀ gas and compared that with galaxy disk rotation and the kinematics of low-ionization-state gas traced by SiĀ IIĀ and SiĀ III. While individual velocity components of the OĀ VIĀ gas do not seem to correlate with the galaxy disk rotation, the column-density-weighted velocity of OĀ VIĀ suggests that the bulk of OĀ VIĀ gas rarely counter-rotate with the disk (FigureĀ 5). We have also matched velocity components of OĀ VIĀ with SiĀ IIĀ and SiĀ IIIĀ by minimizing the differences of their velocity centroids. This analysis has shown that the SiĀ II- or SiĀ III-matched OĀ VIĀ velocity components more frequently corotate with the disk compared to velocity components that cannot be matched with the low ions (FigureĀ 9). In this Section, we discuss the implications and interpretation of these results.
4.1 Location of the Multiphase Gas Traced by OĀ VI, SiĀ II, and SiĀ III
Among the 12 sightlines with OĀ VIĀ detection and spectral coverage of SiĀ IIĀ and SiĀ III, OĀ VIĀ is detected in sightlines out to larger impact parameters compared to the low ions (FiguresĀ 5, 6, and 8). This result is commonly found in previous work. While low ions are only detected in the inner CGM out to kpc around star-forming galaxies, OĀ VIĀ is detected out to impact parameters of several hundred kpcs (see SectionĀ 1). This implies that OĀ VIĀ is more spatially extended compared to the low-ionization-state gas that resides in the inner CGM.
In general, quasar absorption-line analysis only produces line-of-sight measurements and do not reveal the location of the absorbing gas in 3D space. However, our kinematics comparison between OĀ VIĀ and low ions provides the clues for distinguishing the location of the detected gas. FigureĀ 9 shows that SiĀ II- and SiĀ III-matched OĀ VIĀ velocity components more frequently corotate with the disk compared to those components without low-ion matches, and most of the matched components are found in sightlines with impact parameters below 100 kpc. This may not seem surprising, because low ions are often detected in the inner CGM and corotate with the disk, suggesting that the low ions trace the gas residing near the extended disk plane of the galaxy (e.g., Kacprzak etĀ al., 2011; Ho etĀ al., 2017). But with the subset of OĀ VIĀ components having velocities that matched with the low ions and also corotating with the disk, this suggests that these low-ion-matched OĀ VIĀ components trace the gas that is co-spatial with the low ions, i.e., near the extended disk plane in the inner CGM. On the other hand, the unmatched OĀ VIĀ velocity components correspond to gas at the outer CGM.
The interpretation of the location of the absorbing gas explains two results. First, it explains why the SiĀ II- and SiĀ III-matched OĀ VIĀ velocity components are more likely to be corotating with the disk compared to those OĀ VIĀ components without a low-ion match. It also explains the complicated line structures detected in sightlines with small impact parameters (FigureĀ 8), because these sightlines intersect both the inner and outer CGM with different OĀ VIĀ kinematics. Therefore, our results imply that while the outer OĀ VIĀ CGM tends to move randomly, the inner OĀ VIĀ CGM carries angular momentum and corotates with the galaxy disk similar to the low-ionization-state gas.
Furthermore, the OĀ VIĀ velocity components with SiĀ IIĀ matches tend to have larger Doppler parameters compared to those without SiĀ IIĀ matches. FigureĀ 10 shows the distributions of the Doppler parameter and column density for the two OĀ VIĀ populations. Performing the Kolmogorov-Smirnov (KS) test on the Doppler parameter rules out the null hypothesis at 96% that the OĀ VIĀ components with and without SiĀ IIĀ matches are drawn from the same population distribution. However, the null hypothesis cannot be rejected for the column density populations. On the other hand, when we repeat the same analysis for the OĀ VIĀ components with and without SiĀ IIIĀ matches, the null hypothsis cannot be ruled out for both measured quantities. In fact, since SiĀ II-matched OĀ VIĀ components are more likely to be corotating with the disk compared to those without SiĀ IIĀ matches, the larger Doppler parameter of the former plausibly suggests that the rotation motion (in addition to turbulence) has broadened the absorption lines. Because the net CGM angular momentum remains roughly constant with radius in 3D space, the rotation velocity is expected to decrease towards the outer CGM. With SiĀ IIIĀ generally being more spatially extended compared to the lower-ionization SiĀ IIĀ counterpart, we expect the OĀ VIĀ components with SiĀ IIIĀ matches to trace the gas out to larger 3D radii and have lower rotation velocities on average compared to those with SiĀ IIĀ matches. As a result, the rotation motion has produced a measurable increase to the absorption-line width of the OĀ VIĀ velocity components with SiĀ IIĀ matches but not those with SiĀ IIIĀ matches, which also explains why the latter have comparable Doppler parameters as the no-SiĀ III-matched OĀ VIĀ components.
4.2 Angular Momentum of the CGM and Radial Infall
While the absorption-line measurements do not reveal where the gas is located along the line of sight, we have argued that some of the OĀ VIĀ velocity components are tracing gas near the extended disk plane of galaxies (SectionĀ 4.1). But does the corotating gas have sufficient angular momentum to stay on circular orbits? If not, the gas could be infalling to the galaxy, a plausible scenario that explains the corotating low-ionization-state gas at sub-centrifugal velocities (e.g., Ho etĀ al., 2017; Ho & Martin, 2020).
We explore this radial inflow hypothesis by assuming all the non-counter-rotating gas being on the extended disk plane. Then, we calculate the implied galactocentric radii and the circular velocities for individual SiĀ II, SiĀ III, and OĀ VIĀ velocity components. We compare these circular velocities to that required for the gas to stay on circular orbits, assuming the galaxy disks maintain the flat, arctangent rotation curve out to the galactocentric radii. The histograms in FigureĀ 11 show the fractions of SiĀ II, SiĀ III, and OĀ VIĀ velocity components with circular velocities too low to be on circular orbits, i.e., the gas lacks sufficient angular momentum and is infalling (filled histograms)131313 Velocity components that are consistent with being at the galaxy systemic velocities are considered as lacking sufficient angular momentum to stay on circular orbits. . The fraction of infalling velocity components clearly increases with decreasing ionization potential, for which only about 45% for OĀ VIĀ velocity components are tracing infalling gas, and this fraction increases to 80% for SiĀ II. On the other hand, about 15% of the velocity components of each ion have excess angular momentum to be on the circular orbits (hatched). This gas is not expected to be accreting onto the disk.
With the infall interpretation, we can also determine the radii where the infalling gas joins the disk. Assuming the specific angular momentum of the gas remains unchanged, i.e., stays constant, we calculate the radius where the gas has the same specific angular momentum as the extended disk (also known as the circularization radius). For SiĀ II, SiĀ III, and OĀ VIĀ gas, the calculation produces typical radii of under 50 kpc, which is comparable to the radii of extraplanar HĀ IĀ gas of nearby disk galaxies from 21-cm emission observations (e.g., Marasco etĀ al., 2019).
The caveat of the above calculations is that while we assume the angular momentum is conserved for the gas clouds associated with individual velocity components, this may not be case according to cosmological simulations. Simulations have shown that the CGM has a broad angular momentum distribution (Stewart etĀ al., 2013; DeFelippis etĀ al., 2020; Hafen etĀ al., 2022). This is due to the large turbulent velocity of the CGM, for which the turbulent motion adds a random angular momentum vector to the gas clouds (Fielding etĀ al., 2017; Lochhaas etĀ al., 2020; Pandya etĀ al., 2023; Kakoly etĀ al., 2025). However, this turbulent angular momentum component tends to cancel out when the gas accretes onto the galaxy disk (Hafen etĀ al., 2022). In other words, the gas clouds often interact and exchange angular momentum with each other, implying that the angular momentum of a particular gas cloud is unlikely to be conserved during its infall to the disk.
Therefore, we repeat the analysis by comparing the circular velocities required for the gas to stay on circular orbits to the column-density-weighted velocities of different ions in individual sightlines. This enables us to consider the average angular momentum of the gas intersected along the sightlines instead of the gas clouds associated with individual velocity components. The result is shown in FigureĀ 12. First of all, the fraction of counter-rotating OĀ VIĀ gas is only about 10%, which is four times lower than that calculated using individual velocity components (%). This reiterates the result from FigureĀ 5 that net counter-rotating OĀ VIĀ absorbers are rare. Also, the infalling gas fractions of SiĀ II, SiĀ III, and OĀ VIĀ gas all reach %, which are higher than the calculations with individual velocity components, especially for OĀ VIĀ (45%). Because OĀ VIĀ traces gas out to larger 3D radii compared to the low ions, and the expected CGM rotation velocity decreases with increasing radius due to the net CGM angular momentum being roughly radius-independent (also see SectionĀ 4.1), the OĀ VIĀ corotation and/or infall signatures could be easily masked by the turbulent motion. Consequently, the corotation and/or infall of OĀ VIĀ gas is more likely to be observed only when we consider the average angular momentum using the column-density-weighted velocities in individual sightlines instead of individual velocity components.
4.3 Comparison with previous work on OĀ VIĀ kinematics and disk rotation
Unlike the numerous literatures that analyzed the relationship between low-ionization-state MgĀ IIĀ gas and disk rotation, very few work focused on this comparison with OĀ VIĀ (see SectionĀ 1). Using 10 major-axis sightlines with azimuthal angle below 25ā, Kacprzak etĀ al. (2019) showed that both the individual velocity components and weighted velocities per sightline of the OĀ VIĀ gas does not correlate with the disk rotation kinematics (their FigureĀ 4). They also demonstrated that a simple monolithic rotating halo model failed to explain the full velocity spread of the OĀ VIĀ absorption. In contrast, when they ran AMR cosmological simulations, they clearly found infalling OĀ VIĀ filaments that rotate in the same direction as the galaxy disks. The inflowing OĀ VIĀ gas decelerates when it approaches the galaxies and becomes rotationally-dominated. However, the inflowing/rotating OĀ VIĀ signatures were washed out by other velocity structures in the CGM when they computed integrated velocities along lines of sight through the entire simulated galaxy halos. Therefore, inspired by their simulation analysis, they concluded that while corotating or accreting OĀ VIĀ gas could still exist in their observed sightlines, the signature was either weak or was masked by other kinematics ongoing within the CGM.
More recently, Kacprzak etĀ al. (2025) conducted the COS-EDGES survey and analyzed the multiphase CGM of nine edge-on disk galaxies (ā) probed by major-axis sightlines at impact parameters below 0.3. By comparing the stacked absorption-line profiles of different ions (e.g., MgĀ I, MgĀ II, and OĀ VI) in the small () and large () impact parameter bins, they showed that while the Doppler sign of the weighted velocities are consistent with disk rotation, the velocity magnitude decreases with increasing impact parameters. More importantly, the amount of absorption that corotates with the disk declines with both increasing ionization potential and sightline impact parameters; in contrast to over 90% of the MgĀ IIĀ absorption that corotates with the disk, only 80% (60%) of the OĀ VIĀ absorption corotates at small (large) impact parameters. Hence, they concluded that the CGM has a radially-dependent kinematic structure. Within 0.2, the CGM hosts cool and dynamically broad gas that tightly coupled with disk rotation, whereas beyond 0.2, especially for OĀ VI, the CGM becomes less kinematically correlated with disk rotation due to the presence of corotating, lagging, and volume-filling components.
In general, our analysis agrees with both Kacprzak etĀ al. (2019) and Kacprzak etĀ al. (2025) that the corotation signature of OĀ VIĀ gas is challenging to observe due to kinematics components other than rotation. In particular, while individual OĀ VIĀ velocity components do not seem to correlate with disk rotation, the column-density-weighted velocities suggest that the bulk of OĀ VIĀ gas rarely counter-rotates (FigureĀ 5). We have attributed this ādiscrepancyā to the turbulent motion of the CGM, and averaging is required to reveal the net OĀ VIĀ CGM angular momentum while searching for the corotation signature (SectionĀ 4.2). Furthermore, we have also compared the OĀ VIĀ kinematics to that of the low ions. Not only do low ions more often corotate with the disk compared to OĀ VIĀ (FigureĀ 6), but the OĀ VIĀ velocity components with low-ion matches are more likely to corotate compared to those without low-ion matches (FigureĀ 8). These results suggest that low ions are better indicators for identifying the corotating gas, which also matches the conclusion from Nateghi etĀ al. (2024) and Kacprzak etĀ al. (2025) that the lower the ionization potential, the more likely the ion is kinematically consistent with the rotation of the galaxy disks and supports the accretion scenario.
5 Conclusion
We presented a compilation of new and archival HST/COS observations of quasars behind , star-forming galaxies. All 18 quasar sightlines probe the CGM within 45āĀ of the major axes of the host galaxies with inclined disks (ā), doubling the existing number of major-axis sightlines in the literature for studying the relationship between the OĀ VIĀ and disk rotation kinematics.
We compared the kinematics of the OĀ VIĀ gas with that of the low-ionization-state gas (traced by SiĀ IIĀ and SiĀ III) and disk rotation. We showed that while individual velocity OĀ VIĀ components do not correlate with disk rotation, the bulk of OĀ VIĀ gas, as measured by the column-density-weighted OĀ VIĀ velocity in each sightline, rarely counter-rotates with the galaxy disk (FigureĀ 5). In general, SiĀ IIĀ and SiĀ IIIĀ gas are less likely to counter-rotate with the disk compared to OĀ VIĀ (FiguresĀ 6 and 7). But more importantly, we matched velocity components between OĀ VIĀ and low ions (SiĀ IIĀ and SiĀ III) by minimizing the difference of their velocity centroids. We showed that the OĀ VIĀ velocity components with matches with low ions were mainly found in sightlines at small impact parameters of below 100 kpc, whereas those OĀ VIĀ velocity components without low-ion matches were on average at larger impact parameters. These low-ion-matched OĀ VIĀ velocity components are also more frequently corotate with the disk compared to those OĀ VIĀ velocity components without a low-ion match (FigureĀ 9).
Previous literature has shown that low-ionization-state gas is commonly detected in the inner CGM, and major-axis sightlines often detect such gas to corotate with the galaxy disk. Although absorption-line measurements do not reveal where the gas lies along the line of sight, our results of the low-ion matched OĀ VIĀ velocity components sharing these properties suggest that these OĀ VIĀ components trace the gas that is co-spatial with that of the low ions, i.e., on the extended disk plane in the inner CGM. In contrast, the OĀ VIĀ velocity components without a low-ion match traces gas at large 3D radius. Furthermore, if some of the OĀ VIĀ gas resides on the extended disk plane, we asked the question that whether the gas has sufficient angular momentum to stay on circular orbits. If not, this could suggest the gas is infalling. We showed that about 45% of the OĀ VIĀ velocity components lack sufficient angular momentum to maintain a circular orbit. This fraction increases with decreasing ionization potential, with 80% of SiĀ IIĀ velocity components tracing radially infalling gas (FigureĀ 11). Interestingly, repeating this calculation using the column-density-weighted velocities increases the inflow fractions of OĀ VIĀ and SiĀ IIIĀ also to 80%. We interpret this result as a consequence of the turbulent motion adding random angular momentum vectors to the gas clouds associated with individual velocity components. Therefore, the corotation and inflow signatures are more likely to be observationally detected only when we consider the average angular momentum of the gas using the column-density-weighted velocities. Our future work will include photoionization modeling and geometric models for explaining the circumgalactic gas flows of OĀ VIĀ and low-ionization-state gas and to understand the CGM angular momentum and gas accretion onto galaxies.
This work was supported by NASA through HST-GO-15866.002-A from the Space Telescope Science Institute. We gratefully acknowledge the support for SHH from the National Science Foundation under AST-2202328. JS was supported by the Israel Science Foundation (grant No.Ā 2584/21). Some observations reported here were obtained at the MMT Observatory, a joint facility of the University of Arizona and the Smithsonian Institution. Some of the data presented herein were obtained at the W.Ā M.Ā Keck Observatory, which is operated as a scientific partnership among the California Institute of Technology, the University of California and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. Some observations were supported by Swinburne Keck programmes 2014B W018E, 2015 W187E, and 2016A W032E, W056E. Some presented data are based on observations obtained with the Apache Point Observatory 3.5 m telescope, which is owned and operated by the Astrophysical Research Consortium.
Appendix A Absorption System Associated with Individual Galaxies
This Appendix includes the Voigt profile fits of absorption associated with the target galaxies in individual sightlines. FigureĀ 13 shows the Voigt profile fits, and TableĀ 6 lists the parameters of the fits. Because not all ions were used in the analysis of this paper, we only present the measurements in the Appendix.
Quasar | Galaxy | ion | ||||
(km s-1) | (km s-1) | () | ||||
J040148054056 | J040150054047 | 0.21966 | HĀ I | |||
J040148054056 | J040150054047 | 0.21966 | HĀ I | |||
J040148054056 | J040150054047 | 0.21966 | HĀ I | |||
J040148054056 | J040150054047 | 0.21966 | HĀ I | |||
J040148054056 | J040150054047 | 0.21966 | OĀ VI | |||
J040148054056 | J040150054047 | 0.21966 | OĀ VI | |||
J040148054056 | J040150054047 | 0.21966 | CĀ III | |||
J040148054056 | J040150054047 | 0.21966 | CĀ III | |||
J040148054056 | J040150054047 | 0.21966 | CĀ III | |||
J040148054056 | J040150054047 | 0.21966 | NĀ III | |||
J040148054056 | J040150054047 | 0.21966 | NĀ III | |||
J040148054056 | J040150054047 | 0.21966 | SiĀ III | |||
J040148054056 | J040150054047 | 0.21966 | SiĀ III | |||
J040148054056 | J040150054047 | 0.21966 | SiĀ IV | |||
J040148054056 | J040150054047 | 0.21966 | SĀ IV | |||
J080704+360353 | J080702+360141 | 0.088061 | HĀ I | |||
J080704+360353 | J080702+360141 | 0.088061 | HĀ I | |||
J080704+360353 | J080702+360141 | 0.088061 | HĀ I | |||
J080704+360353 | J080702+360141 | 0.088061 | HĀ I | |||
J080704+360353 | J080702+360141 | 0.088061 | HĀ I | |||
J080704+360353 | J080702+360141 | 0.088061 | HĀ I | |||
J080704+360353 | J080702+360141 | 0.088061 | HĀ I | |||
J080704+360353 | J080702+360141 | 0.088061 | HĀ I | |||
J080704+360353 | J080702+360141 | 0.088061 | OĀ VI | |||
J080704+360353 | J080702+360141 | 0.088061 | OĀ VI | |||
J080704+360353 | J080702+360141 | 0.088061 | OĀ VI | |||
J080704+360353 | J080702+360141 | 0.088061 | CĀ IV | |||
J080704+360353 | J080702+360141 | 0.088061 | CĀ IV | |||
J080704+360353 | J080702+360141 | 0.088061 | CĀ IV | |||
J080704+360353 | J080702+360141 | 0.088061 | NĀ V | |||
J084349+261910 | J084356+261855 | 0.11284 | HĀ I | |||
J084349+261910 | J084356+261855 | 0.11284 | HĀ I | |||
J084349+261910 | J084356+261855 | 0.11284 | HĀ I | |||
J084349+261910 | J084356+261855 | 0.11284 | OĀ VI | |||
J084349+261910 | J084356+261855 | 0.11284 | OĀ VI | |||
J084349+261910 | J084356+261855 | 0.11284 | SiĀ III | |||
J084349+261910 | J084356+261855 | 0.11284 | SiĀ III | |||
J084349+261910 | J084356+261855 | 0.11284 | SiĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | HĀ I | |||
J085215+171143 | J085215+171137 | 0.16921 | HĀ I | |||
J085215+171143 | J085215+171137 | 0.16921 | HĀ I | |||
J085215+171143 | J085215+171137 | 0.16921 | HĀ I | |||
J085215+171143 | J085215+171137 | 0.16921 | CĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | CĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | OĀ VI | |||
J085215+171143 | J085215+171137 | 0.16921 | OĀ VI | |||
J085215+171143 | J085215+171137 | 0.16921 | OĀ VI | |||
J085215+171143 | J085215+171137 | 0.16921 | OĀ VI | |||
J085215+171143 | J085215+171137 | 0.16921 | CĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | CĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | CĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | CĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | NĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | NĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | NĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | NĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | NĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | NĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | NĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | NĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | NĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | OĀ I | |||
J085215+171143 | J085215+171137 | 0.16921 | OĀ I | |||
J085215+171143 | J085215+171137 | 0.16921 | OĀ I | |||
J085215+171143 | J085215+171137 | 0.16921 | OĀ I | |||
J085215+171143 | J085215+171137 | 0.16921 | OĀ I | |||
J085215+171143 | J085215+171137 | 0.16921 | OĀ I | |||
J085215+171143 | J085215+171137 | 0.16921 | SĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | SĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | SĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | SiĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | SiĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | SiĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | SiĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | SiĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | SiĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | SiĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | FeĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | FeĀ II | |||
J085215+171143 | J085215+171137 | 0.16921 | FeĀ III | |||
J085215+171143 | J085215+171137 | 0.16921 | FeĀ III | |||
J091954+291408 | J091954+291345 | 0.23288 | HĀ I | |||
J091954+291408 | J091954+291345 | 0.23288 | HĀ I | |||
J091954+291408 | J091954+291345 | 0.23288 | HĀ I | |||
J091954+291408 | J091954+291345 | 0.23288 | HĀ I | |||
J091954+291408 | J091954+291345 | 0.23288 | CĀ II | |||
J091954+291408 | J091954+291345 | 0.23288 | CĀ II | |||
J091954+291408 | J091954+291345 | 0.23288 | CĀ II | |||
J091954+291408 | J091954+291345 | 0.23288 | OĀ VI | |||
J091954+291408 | J091954+291345 | 0.23288 | OĀ VI | |||
J091954+291408 | J091954+291345 | 0.23288 | OĀ VI | |||
J091954+291408 | J091954+291345 | 0.23288 | OĀ VI | |||
J091954+291408 | J091954+291345 | 0.23288 | OĀ VI | |||
J091954+291408 | J091954+291345 | 0.23288 | OĀ VI | |||
J091954+291408 | J091954+291345 | 0.23288 | CĀ III | |||
J091954+291408 | J091954+291345 | 0.23288 | CĀ III | |||
J091954+291408 | J091954+291345 | 0.23288 | CĀ III | |||
J091954+291408 | J091954+291345 | 0.23288 | CĀ III | |||
J091954+291408 | J091954+291345 | 0.23288 | CĀ III | |||
J091954+291408 | J091954+291345 | 0.23288 | NĀ II | |||
J091954+291408 | J091954+291345 | 0.23288 | NĀ II | |||
J094331+053131 | J094330+053118 | 0.353052 | HĀ I | |||
J094331+053131 | J094330+053118 | 0.353052 | HĀ I | |||
J094331+053131 | J094330+053118 | 0.353052 | HĀ I | |||
J094331+053131 | J094330+053118 | 0.353052 | HĀ I | |||
J094331+053131 | J094330+053118 | 0.353052 | HĀ I | |||
J094331+053131 | J094330+053118 | 0.353052 | HĀ I | |||
J094331+053131 | J094330+053118 | 0.353052 | HĀ I | |||
J094331+053131 | J094330+053118 | 0.353052 | HĀ I | |||
J094331+053131 | J094330+053118 | 0.353052 | OĀ VI | |||
J094331+053131 | J094330+053118 | 0.353052 | OĀ VI | |||
J094331+053131 | J094330+053118 | 0.353052 | OĀ VI | |||
J094331+053131 | J094330+053118 | 0.353052 | CĀ III | |||
J094331+053131 | J094330+053118 | 0.353052 | CĀ III | |||
J094331+053131 | J094330+053118 | 0.353052 | CĀ III | |||
J094331+053131 | J094330+053118 | 0.353052 | CĀ III | |||
J094331+053131 | J094330+053118 | 0.353052 | CĀ III | |||
J094331+053131 | J094330+053118 | 0.353052 | CĀ III | |||
J094331+053131 | J094330+053118 | 0.353052 | CĀ III | |||
J094331+053131 | J094330+053118 | 0.353052 | CĀ III | |||
J094331+053131 | J094330+053118 | 0.353052 | CĀ III | |||
J094331+053131 | J094330+053118 | 0.353052 | NĀ V | |||
J094331+053131 | J094330+053118 | 0.353052 | NĀ V | |||
J094331+053131 | J094330+053118 | 0.353052 | NĀ V | |||
J094331+053131 | J094330+053118 | 0.353052 | SiĀ III | |||
J103640+565125 | J103643+565119 | 0.13629 | HĀ I | |||
J103640+565125 | J103643+565119 | 0.13629 | HĀ I | |||
J103640+565125 | J103643+565119 | 0.13629 | SiĀ III | |||
J103640+565125 | J103643+565119 | 0.13629 | SiĀ III | |||
J104116+061016 | J104117+061018 | 0.442173 | HĀ I | |||
J104116+061016 | J104117+061018 | 0.442173 | HĀ I | |||
J104116+061016 | J104117+061018 | 0.442173 | HĀ I | |||
J104116+061016 | J104117+061018 | 0.442173 | HĀ I | |||
J104116+061016 | J104117+061018 | 0.442173 | HĀ I | |||
J104116+061016 | J104117+061018 | 0.442173 | CĀ II | |||
J104116+061016 | J104117+061018 | 0.442173 | CĀ II | |||
J104116+061016 | J104117+061018 | 0.442173 | CĀ II | |||
J104116+061016 | J104117+061018 | 0.442173 | CĀ II | |||
J104116+061016 | J104117+061018 | 0.442173 | OĀ VI | |||
J104116+061016 | J104117+061018 | 0.442173 | OĀ VI | |||
J104116+061016 | J104117+061018 | 0.442173 | OĀ VI | |||
J104116+061016 | J104117+061018 | 0.442173 | CĀ III | |||
J104116+061016 | J104117+061018 | 0.442173 | CĀ III | |||
J104116+061016 | J104117+061018 | 0.442173 | NĀ II | |||
J104116+061016 | J104117+061018 | 0.442173 | NĀ III | |||
J104116+061016 | J104117+061018 | 0.442173 | NĀ III | |||
J104116+061016 | J104117+061018 | 0.442173 | SiĀ II | |||
J104116+061016 | J104117+061018 | 0.442173 | SiĀ II | |||
J104116+061016 | J104117+061018 | 0.442173 | SiĀ II | |||
J104116+061016 | J104117+061018 | 0.442173 | SiĀ II | |||
J104116+061016 | J104117+061018 | 0.442173 | SiĀ II | |||
J104116+061016 | J104117+061018 | 0.442173 | SiĀ III | |||
J104116+061016 | J104117+061018 | 0.442173 | SiĀ III | |||
J104116+061016 | J104117+061018 | 0.442173 | SiĀ III | |||
J104116+061016 | J104117+061018 | 0.442173 | SiĀ III | |||
J113910135043 | J113911135108 | 0.204194 | HĀ I | |||
J113910135043 | J113911135108 | 0.204194 | HĀ I | |||
J113910135043 | J113911135108 | 0.204194 | HĀ I | |||
J113910135043 | J113911135108 | 0.204194 | HĀ I | |||
J113910135043 | J113911135108 | 0.204194 | HĀ I | |||
J113910135043 | J113911135108 | 0.204194 | HĀ I | |||
J113910135043 | J113911135108 | 0.204194 | HĀ I | |||
J113910135043 | J113911135108 | 0.204194 | HĀ I | |||
J113910135043 | J113911135108 | 0.204194 | HĀ I | |||
J113910135043 | J113911135108 | 0.204194 | CĀ II | |||
J113910135043 | J113911135108 | 0.204194 | CĀ II | |||
J113910135043 | J113911135108 | 0.204194 | OĀ VI | |||
J113910135043 | J113911135108 | 0.204194 | OĀ VI | |||
J113910135043 | J113911135108 | 0.204194 | OĀ VI | |||
J113910135043 | J113911135108 | 0.204194 | CĀ III | |||
J113910135043 | J113911135108 | 0.204194 | CĀ III | |||
J113910135043 | J113911135108 | 0.204194 | CĀ III | |||
J113910135043 | J113911135108 | 0.204194 | CĀ III | |||
J113910135043 | J113911135108 | 0.204194 | CĀ III | |||
J113910135043 | J113911135108 | 0.204194 | NĀ III | |||
J113910135043 | J113911135108 | 0.204194 | NĀ III | |||
J113910135043 | J113911135108 | 0.204194 | SiĀ III | |||
J113910135043 | J113911135108 | 0.204194 | SiĀ III | |||
J113910135043 | J113911135108 | 0.204194 | SiĀ III | |||
J113910135043 | J113911135108 | 0.204194 | SiĀ III | |||
J113910135043 | J113911135108 | 0.204194 | SiĀ III | |||
J113910135043 | J113911135108 | 0.204194 | SiĀ III | |||
J113910135043 | J113909135053 | 0.319255 | HĀ I | |||
J113910135043 | J113909135053 | 0.319255 | HĀ I | |||
J113910135043 | J113909135053 | 0.319255 | HĀ I | |||
J113910135043 | J113909135053 | 0.319255 | OĀ VI | |||
J113910135043 | J113909135053 | 0.319255 | OĀ VI | |||
J113910135043 | J113909135053 | 0.319255 | OĀ VI | |||
J113910135043 | J113909135053 | 0.319255 | CĀ III | |||
J113910135043 | J113909135053 | 0.319255 | CĀ III | |||
J113910135043 | J113909135053 | 0.319255 | CĀ III | |||
J113910135043 | J113909135053 | 0.319255 | CĀ III | |||
J113910135043 | J113909135053 | 0.319255 | CĀ III | |||
J123335+475800 | J123338+475757 | 0.22171 | HĀ I | |||
J123335+475800 | J123338+475757 | 0.22171 | HĀ I | |||
J123335+475800 | J123338+475757 | 0.22171 | HĀ I | |||
J123335+475800 | J123338+475757 | 0.22171 | HĀ I | |||
J123335+475800 | J123338+475757 | 0.22171 | CĀ II | |||
J123335+475800 | J123338+475757 | 0.22171 | CĀ II | |||
J123335+475800 | J123338+475757 | 0.22171 | CĀ II | |||
J123335+475800 | J123338+475757 | 0.22171 | OĀ VI | |||
J123335+475800 | J123338+475757 | 0.22171 | CĀ III | |||
J123335+475800 | J123338+475757 | 0.22171 | CĀ III | |||
J123335+475800 | J123338+475757 | 0.22171 | CĀ III | |||
J123335+475800 | J123338+475757 | 0.22171 | NĀ II | |||
J123335+475800 | J123338+475757 | 0.22171 | NĀ II | |||
J123335+475800 | J123338+475757 | 0.22171 | NĀ III | |||
J123335+475800 | J123338+475757 | 0.22171 | NĀ III | |||
J123335+475800 | J123338+475757 | 0.22171 | SiĀ II | |||
J123335+475800 | J123338+475757 | 0.22171 | SiĀ II | |||
J123335+475800 | J123338+475757 | 0.22171 | SiĀ III | |||
J123335+475800 | J123338+475757 | 0.22171 | SiĀ III | |||
J123335+475800 | J123338+475757 | 0.22171 | SiĀ III | |||
J123335+475800 | J123338+475757 | 0.22171 | SiĀ IV | |||
J132222+464546 | J132222+464546 | 0.214431 | HĀ I | |||
J132222+464546 | J132222+464546 | 0.214431 | HĀ I | |||
J132222+464546 | J132222+464546 | 0.214431 | CĀ II | |||
J132222+464546 | J132222+464546 | 0.214431 | CĀ II | |||
J132222+464546 | J132222+464546 | 0.214431 | OĀ VI | |||
J132222+464546 | J132222+464546 | 0.214431 | OĀ VI | |||
J132222+464546 | J132222+464546 | 0.214431 | CĀ III | |||
J132222+464546 | J132222+464546 | 0.214431 | CĀ III | |||
J132222+464546 | J132222+464546 | 0.214431 | NĀ II | |||
J132222+464546 | J132222+464546 | 0.214431 | NĀ III | |||
J132222+464546 | J132222+464546 | 0.214431 | NĀ III | |||
J132222+464546 | J132222+464546 | 0.214431 | NĀ V | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ II | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ II | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ II | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ II | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ III | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ III | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ III | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ III | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ III | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ IV | |||
J132222+464546 | J132222+464546 | 0.214431 | SiĀ IV | |||
J135522+303324 | J135521+303320 | 0.20690 | HĀ I | |||
J135522+303324 | J135521+303320 | 0.20690 | HĀ I | |||
J135522+303324 | J135521+303320 | 0.20690 | CĀ II | |||
J135522+303324 | J135521+303320 | 0.20690 | OĀ VI | |||
J135522+303324 | J135521+303320 | 0.20690 | OĀ VI | |||
J135522+303324 | J135521+303320 | 0.20690 | CĀ III | |||
J135522+303324 | J135521+303320 | 0.20690 | NĀ III | |||
J135522+303324 | J135521+303320 | 0.20690 | NĀ III | |||
J142501+382100 | J142459+382113 | 0.21295 | HĀ I | |||
J142501+382100 | J142459+382113 | 0.21295 | HĀ I | |||
J142501+382100 | J142459+382113 | 0.21295 | HĀ I | |||
J142501+382100 | J142459+382113 | 0.21295 | OĀ VI | |||
J142501+382100 | J142459+382113 | 0.21295 | OĀ VI | |||
J142501+382100 | J142459+382113 | 0.21295 | CĀ III | |||
J142501+382100 | J142459+382113 | 0.21295 | CĀ III | |||
J142501+382100 | J142459+382113 | 0.21295 | NĀ II | |||
J154741+343357 | J154741+343350 | 0.18392 | HĀ I | |||
J154741+343357 | J154741+343350 | 0.18392 | HĀ I | |||
J154741+343357 | J154741+343350 | 0.18392 | OĀ VI | |||
J154741+343357 | J154741+343350 | 0.18392 | OĀ VI | |||
J154741+343357 | J154741+343350 | 0.18392 | OĀ VI | |||
J154741+343357 | J154741+343350 | 0.18392 | OĀ VI | |||
J154741+343357 | J154741+343350 | 0.18392 | CĀ II | |||
J154741+343357 | J154741+343350 | 0.18392 | CĀ II | |||
J154741+343357 | J154741+343350 | 0.18392 | CĀ III | |||
J154741+343357 | J154741+343350 | 0.18392 | CĀ III | |||
J154741+343357 | J154741+343350 | 0.18392 | NĀ III | |||
J154741+343357 | J154741+343350 | 0.18392 | NĀ III | |||
J154741+343357 | J154741+343350 | 0.18392 | SiĀ II | |||
J154741+343357 | J154741+343350 | 0.18392 | SiĀ II | |||
J154741+343357 | J154741+343350 | 0.18392 | SiĀ II | |||
J154741+343357 | J154741+343350 | 0.18392 | SiĀ II | |||
J154741+343357 | J154741+343350 | 0.18392 | SiĀ III | |||
J154741+343357 | J154741+343350 | 0.18392 | SiĀ III | |||
J154741+343357 | J154741+343350 | 0.18392 | FeĀ III | |||
J154741+343357 | J154741+343350 | 0.18392 | FeĀ III | |||
J154741+343357 | J154741+343350 | 0.18392 | FeĀ III | |||
J155504+362847 | J155505+362848 | 0.18926 | HĀ I | |||
J155504+362847 | J155505+362848 | 0.18926 | CĀ II | |||
J155504+362847 | J155505+362848 | 0.18926 | CĀ II | |||
J155504+362847 | J155505+362848 | 0.18926 | OĀ VI | |||
J155504+362847 | J155505+362848 | 0.18926 | OĀ VI | |||
J155504+362847 | J155505+362848 | 0.18926 | CĀ III | |||
J155504+362847 | J155505+362848 | 0.18926 | NĀ II | |||
J155504+362847 | J155505+362848 | 0.18926 | NĀ III | |||
J155504+362847 | J155505+362848 | 0.18926 | NĀ III | |||
J155504+362847 | J155505+362848 | 0.18926 | SiĀ II | |||
J155504+362847 | J155505+362848 | 0.18926 | SiĀ II | |||
J155504+362847 | J155505+362848 | 0.18926 | SiĀ III | |||
J155504+362847 | J155505+362848 | 0.18926 | SiĀ III | |||
J155504+362847 | J155505+362848 | 0.18926 | SiĀ IV | |||
J155504+362847 | J155505+362848 | 0.18926 | SiĀ IV | |||
J160951+353843 | J160951+353838 | 0.28940 | HĀ I | |||
J160951+353843 | J160951+353838 | 0.28940 | OĀ I | |||
J160951+353843 | J160951+353838 | 0.28940 | OĀ I | |||
J160951+353843 | J160951+353838 | 0.28940 | OĀ I | |||
J160951+353843 | J160951+353838 | 0.28940 | CĀ II | |||
J160951+353843 | J160951+353838 | 0.28940 | CĀ II | |||
J160951+353843 | J160951+353838 | 0.28940 | CĀ II | |||
J160951+353843 | J160951+353838 | 0.28940 | CĀ II | |||
J160951+353843 | J160951+353838 | 0.28940 | OĀ VI | |||
J160951+353843 | J160951+353838 | 0.28940 | OĀ VI | |||
J160951+353843 | J160951+353838 | 0.28940 | OĀ VI | |||
J160951+353843 | J160951+353838 | 0.28940 | OĀ VI | |||
J160951+353843 | J160951+353838 | 0.28940 | OĀ VI | |||
J160951+353843 | J160951+353838 | 0.28940 | CĀ III | |||
J160951+353843 | J160951+353838 | 0.28940 | NĀ II | |||
J160951+353843 | J160951+353838 | 0.28940 | NĀ II | |||
J160951+353843 | J160951+353838 | 0.28940 | NĀ II | |||
J160951+353843 | J160951+353838 | 0.28940 | NĀ II | |||
J160951+353843 | J160951+353838 | 0.28940 | NĀ II | |||
J160951+353843 | J160951+353838 | 0.28940 | SĀ III | |||
J160951+353843 | J160951+353838 | 0.28940 | SĀ IV | |||
J225357+160853 | J225400+160925 | 0.390013 | HĀ I | |||
J225357+160853 | J225400+160925 | 0.390013 | HĀ I | |||
J225357+160853 | J225400+160925 | 0.390013 | HĀ I | |||
J225357+160853 | J225400+160925 | 0.390013 | HĀ I | |||
J225357+160853 | J225400+160925 | 0.390013 | HĀ I | |||
J225357+160853 | J225400+160925 | 0.390013 | HĀ I | |||
J225357+160853 | J225400+160925 | 0.390013 | HĀ I | |||
J225357+160853 | J225400+160925 | 0.390013 | HĀ I | |||
J225357+160853 | J225400+160925 | 0.390013 | HĀ I | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ VI | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ VI | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ VI | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ VI | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ VI | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ VI | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ VI | |||
J225357+160853 | J225400+160925 | 0.390013 | CĀ III | |||
J225357+160853 | J225400+160925 | 0.390013 | CĀ III | |||
J225357+160853 | J225400+160925 | 0.390013 | CĀ III | |||
J225357+160853 | J225400+160925 | 0.390013 | NĀ III | |||
J225357+160853 | J225400+160925 | 0.390013 | NĀ V | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ III | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ III | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ III | |||
J225357+160853 | J225400+160925 | 0.390013 | OĀ III |